Need Help? Raxd it 3. O 0/2 points I s Answers SerCP10 9.P.012 MI FB Assume that if the shear stress in steel exceeds about 4.00 × 108 N/m2, the steel ruptures. (a) Determine the shearing force necessary to shear a steel bolt 1.40 cm in diameter Apply the definition of the shear stress to determine the force needed to shear the steel bolt as described. N b) Determine the shearing force necessary to punch a 1.50-cm-diameter hole in a steel plate 0.580 cm thick Need Help? ReadItMaster t Submit Answer Save Progress Pra Practice Another Version
Answer #1

a) r = d / 2 = 1.4 / 2 = 0.7 cm

force = stress * area

= 4.00 * 108 * pi * (0.7 * 10-2)2= 6.157 * 104 N

b) surface area = pi * 0.015 * 0.580 * 10-2 = 2.73 * 10-4

force = stress * surface area

= 4.00 * 108 * 2.73 * 10-4= 1.09 * 105 N